### v1

@@ Abstract
Achieving high-fidelity 3D surface reconstruction while preserving fine details remains challenging, especially in the presence of materials with complex reflectance properties and without a dense-view setup. In this paper, we introduce a versatile framework that incorporates multi-view normal and optionally reflectance maps into radiance-based surface reconstruction. Our approach employs a pixel-wise joint re-parametrization of reflectance and surface normals, representing them as a vector of radiances under simulated, varying illumination. This formulation enables seamless incorporation into standard surface reconstruction pipelines, such as traditional multi-view stereo (MVS) frameworks or modern neural volume rendering (NVR) ones. Combined with the latter, our approach achieves state-of-the-art performance on multi-view photometric stereo (MVPS) benchmark datasets, including DiLiGenT-MV, LUCES-MV and Skoltech3D. In particular, our method excels in reconstructing fine-grained details and handling challenging visibility conditions. The present paper is an extended version of the earlier conference paper by Brument et al. (in Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), 2024), featuring an accelerated and more robust algorithm as well as a broader empirical evaluation. The code and data relative to this article is available at https://github.com/RobinBruneau/RNb-NeuS2.
@@ 1 Introduction
The recovery of fine-grained structures is a long-standing bottleneck in 3D surface reconstruction. Traditional multi-view stereo (MVS) methods [25, 73] often produce overly smoothed surfaces, struggling with sharp edges and intricate geometries [74, 26]. Recent neural approaches have advanced surface reconstruction in this direction by shifting from local, patch-based optimization to global, pixel-wise optimization [94, 95, 64, 82]. Notably, methods such as HF-NeuS [84], PET-NeuS [86], NeuS2 [85], and Neuralangelo [52] excel in capturing fine surface details – at least, when the density of viewpoints is high [7, 56].
@@
This paper builds upon and significantly extends our previous conference publication [7], where the foundational concepts of this two-stage approach were introduced. Therein, we presented the key idea to re-parametrize normal and reflectance priors into simulated radiance values on a per-pixel basis, facilitating seamless incorporation into existing surface reconstruction pipelines, notably neural volume rendering (NVR) frameworks based on signed distance functions (SDF). This innovative way to perform multi-view normal (and reflectance) integration yields a state-of-the-art multi-view photometric stereo (MVPS) technique, summarized in Figure 1, which we evaluated on the DiLiGenT-MV dataset [51] using the SDM-UniPS method [37] for normal and reflectance estimation.
Figure 1: Overview of the surface reconstruction pipeline “RNb-NeuS” proposed in [7]. From multi-view multi-light data, photometric stereo (noted PS) first estimates per-view normal maps $\{\mathbf{n}_{k}\}$ and optionally reflectance maps $\{\mathbf{r}_{k}\}$. These estimates are re-parametrized into radiance vectors $\mathbf{v}_{k}=\mathsf{F}(\mathbf{r}_{k},\mathbf{n}_{k},\mathsf{L}_{k})$ via a physically-based rendering function $\mathsf{F}$ under simulated lightings $\mathsf{L}_{k}$. An implicit neural representation, consisting of a signed distance function $f$ and a reflectance $\boldsymbol{\rho}$, is then learned by minimising the discrepancy between $\mathbf{v}_{k}$ and its volume-rendered counterpart $\tilde{\mathbf{v}}_{k}(\boldsymbol{\rho},f)$. The final surface mesh is extracted as the zero-level set of $f$ after optimisation.
@@
  • We implement our approach within the NeuS2 framework [85], achieving a 100$\times$ speedup compared to our original implementation (see Section 5.5).
@@
  • In addition to volume rendering, we validate our re-parametrization in a traditional patch-based MVS framework, demonstrating both its broad applicability and its ability to achieve exact integration in a noiseless setup (see Section 4).
@@
The rest of this article is organized as follows. We first review the relevant literature in Section 2. Then, we show in Section 3 how to jointly re-parametrize normal and reflectance data into simulated radiance values. We embed this re-parametrization in a traditional MVS framework in Section 4, to demonstrate the feasibility of exact surface reconstruction from normals and reflectance cues. To ensure robustness, in Section 5 we then turn our attention to replacing this framework by a volume rendering one based on neural implicit surfaces. Extensive experiments validating the proposed approach are then conducted in Section 6, before our conclusions are drawn in Section 7.
@@ 2.1 Normal Estimation
Single-image normal estimation is highly beneficial for 3D scene reconstruction, since it preserves local geometry without metric ambiguity [98, 81]. Recent approaches predominantly leverage learning-based methodologies [49, 83, 67, 68, 54, 19, 4, 92]. However, direct training on real data remains challenging, as normal labels cannot be directly captured by current sensor technology. Typically, normals are thus inferred from depth maps [76, 22], but inaccuracies persist despite various correction strategies [3, 58]. Therefore, the most recent works employed large-scale datasets of synthesized data such as OmniData [21, 42], along with smart architectural choices, e.g. per-pixel ray direction modelling in DSINE [2] or ViTs and diffusion models in StableNormal [96].
Photometric stereo [88], on the other hand, is also a monocular technique yet it analyzes multiple images captured under varying lighting. The historical inverse problem-based PS approach, which fits normals and reflectance to a physics-based image formation model, has been employed in the calibrated (known illumination) scenario [27, 89, 75, 39, 38], in the uncalibrated one [32, 11, 1, 24, 70], as well as under unknown and spatially-varying lighting [5, 69, 61, 30, 29]. However, switching from the inverse problem framework to the deep learning paradigm has recently revolutionized PS. Limitations of traditional physics-based models, especially regarding non-Lambertian effects, have thereby been addressed for both calibrated  [72, 34, 12, 35, 87] and uncalibrated PS [77, 13, 14, 43, 50], and the most modern techniques make lighting context implicit, which both improves performances [41] and allows tackling “universal” (unknown and spatially-varying illumination) PS [36, 37, 31].
In multi-view single-light contexts, traditional methods, such as Colmap’s multi-view stereo [73], infer normal maps by analysing patch orientations [6]. Differently, [9] estimate normals from image grey level variations and use them in a slanted plane-sweeping algorithm. An alternative paradigm computes normals post-MVS by deriving them from depth maps, as demonstrated in RegSDF [99], which utilizes depth from Vis-MVSNet [100]. This depth-based normal estimation is MVS-method agnostic, applicable to both traditional [25, 90] and deep learning-driven methods [93, 28, 100].
@@ 2.2 Reflectance Estimation
In multi-view frameworks, methods such as Ref-NeRF [78] extend neural radiance fields (NeRF) to explicitly model reflections, improving the reconstruction of specular and glossy surfaces. Techniques like SHINOBI [23] concurrently optimize geometry, reflectance, and illumination, enabling detailed relightable 3D assets. More recently, DiffusionRenderer [53] provided a unified solution for high-quality normal and reflectance reconstruction from video sequences, by bridging monocular and multi-view estimation through diffusion models for neural inverse rendering.
@@ 2.3 Multi-view Normal Integration
In a multi-view single-light setting, MonoSDF [98] combined a normal and depth consistency loss derived from monocular predictions [21] with a radiance constraint, a common approach in NeRF-like [60] methods. Similarly, Gaussian Surfels [17] uses monocular normal predictions to constrain 3D Gaussian Splatting (3DGS) [46].
@@
Yet, multi-view normal integration is much more studied under the prism of multi-view photometric stereo. This problem was first addressed by [33], through the optimization of a loss combining a rendering term and a discrepancy between photometric stereo normals and the optimized mesh. Note that this first approach required neither prior knowledge of camera poses nor illumination conditions. Similarly, [65, 66] simultaneously estimated reflectance, normals, and illumination through uncalibrated PS, leveraging the 3D mesh normals to resolve ambiguities and refining surface details. [55] later formulated the problem within an SDF representation, achieving superior surface detail reconstruction compared to [66]. Further refinements were made by [51], who enhanced a 3D mesh by propagating SfM points following the method of [63], and validating their method through the introduction of the publicly available dataset “DiLiGenT-MV”.
@@
Neural surface reconstruction methods have recently emerged as a promising alternative, by enforcing alignment between per-view normal maps and the gradient of a neural SDF. [44] for instance constrained the SDF optimization by CNN-PS normals  [34] and MVS depths [80], incorporating uncertainty measures to mitigate conflicts in predictions. [45] further added a neural volume rendering loss, benefiting the latter’s robustness in handling various material types. This resulted in a multi-objective optimization comprising three loss terms. However, as in [44], the reliance on uncertainty-based hyperparameter tuning did not fully resolve conflicts between objectives, leading to potential loss of fine-grained details. PS-NeRF [91] introduced a two-stage solution, where the first stage achieves multi-view normal integration by aligning surface gradients with SDPS-Net normals [13], and the second stage leverages UNISURF [64] to optimize geometry, reflectance and illumination, modelled by multi-layer perceptrons (MLPs). However, its reliance on the directional illumination assumption limited its generalizability. NPL-MVPS [57] relaxed this assumption by considering a near-light model, initially enforcing alignment between SDF gradients and UniMS-PS normals [31], and subsequently processing to joint shape and reflectance optimization while considering both light attenuation and cast shadows. Most recently, SuperNormal [10] adopted a similar initialization approach to PS-NeRF and NPL-MVPS, utilizing SDM-UniPS normals [37], but significantly improved efficiency with multi-resolution hash encoding and directional finite differences, achieving nearly double the training speed.
@@
In contrast to these prior methods, the novel approach we propose in the rest of this article formulates multi-view normal integration as a single-objective optimization problem, through joint re-parametrization of normals and reflectance. Let us now introduce this re-parametrization.
3 Reflectance and Normal Re-parametrization
@@
Reflectance and surface normals traditionally represent distinct types of information: reflectance captures intrinsic photometric properties of surface materials, while normals describe the geometric orientation of surfaces. This inherent heterogeneity complicates their simultaneous optimization, typically necessitating separate processing steps, e.g., depth triangulation from multi-view data followed by the fusion of this depth information with normal maps using multi-objective optimization frameworks [44, 45].
@@
To address this issue, we introduce a re-parametrization approach designed to unify surface normals and reflectance into homogeneous quantities, by mapping both of them to simulated radiance values under varying illumination conditions. This radiance-based parametrization enables a unified optimization reducing the need for additional regularization, commonly required in multi-objective contexts [98, 99], consequently enhancing consistency and computational efficiency. Indeed, a key motivation for radiance-based re-parametrization is its compatibility with existing photometric cost-minimization frameworks employed in classical multi-view stereo [25, 73] and neural volume rendering methods [82, 95, 52]. This incorporation facilitates optimization of photometric consistency across viewpoints or between input and rendered images.
@@ 3.1 Input Data
Given a set of $N$ viewpoints, we assume the availability of corresponding reflectance maps $\mathsf{R}_{i}$ and normal maps $\mathsf{N}_{i}$, indexed by $i\in\{1,\dots,N\}$. Each viewpoint map comprises $m$ pixels, indexed by $k\in\{1,\dots,m\}$, such that:
$\mathsf{R}_{i}=\{\mathbf{r}_{i,k}\}_{k\in\{1,\dots,m\}},\quad\mathsf{N}_{i}=\{\mathbf{n}_{i,k}\}_{k\in\{1,\dots,m\}}.$ (1)
Therein, the reflectance for the $k$-th pixel in the $i$-th view is given by a parameter vector $\mathbf{r}_{i,k}\in\mathbb{R}^{q}$ with dimension $q$ corresponding to the specific reflectance model used, and the outward normal by a unit vector $\mathbf{n}_{i,k}\in\mathbb{S}^{2}$. The latter is expressed, using the known camera poses, in world coordinates. If reflectance data is unavailable, simply setting $\mathbf{r}_{i,k}=[1]~{}\forall(i,k)$ allows our framework to be used exclusively for multi-view normal integration.
3.2 Re-parametrization
Our re-parametrization approach transforms each couple of reflectance and normal vectors $(\mathbf{r}_{i,k},\mathbf{n}_{i,k})$ into a vector $\mathbf{v}_{i,k}$ of homogeneous radiance values, simulated using a physically-based rendering (PBR) model:
$\mathbf{v}_{i,k}=\mathsf{F}(\mathbf{r}_{i,k},\mathbf{n}_{i,k},\mathsf{L}_{i,k}),$ (2)
with $\mathsf{L}_{i,k}\in\mathbb{R}^{n\times l}$ the illumination conditions (chosen specifically for the $k$-th photosite of the $i$-th view) and $\mathsf{F}:\,\mathbb{R}^{q}\times\mathbb{S}^{2}\times\mathbb{R}^{n\times l}\to\mathbb{R}^{n\times q}$ the PBR function (for simplicity, the same model is used for all pixels). Here, $l$ stands for the dimensionality of the light and $n$ for that of the radiance representation, and both of them depend on the choice of a particular model.

### v3

@@ Abstract
Achieving high-fidelity 3D surface reconstruction while preserving fine details remains challenging, especially in the presence of materials with complex reflectance properties and without a dense-view setup. In this paper, we introduce a versatile framework that incorporates multi-view normal and optionally reflectance maps into radiance-based surface reconstruction. Our approach employs a pixel-wise joint re-parametrisation of reflectance and surface normals, representing them as a vector of radiances under simulated, varying illumination. This formulation enables seamless incorporation into standard surface reconstruction pipelines, such as traditional multi-view stereo (MVS) frameworks or modern neural volume rendering (NVR) ones. Combined with the latter, our approach achieves state-of-the-art performance on multi-view photometric stereo (MVPS) benchmark datasets, including DiLiGenT-MV, LUCES-MV and Skoltech3D. In particular, our method excels in reconstructing fine-grained details and handling challenging visibility conditions. The present paper is an extended version of the earlier conference paper by Brument et al (in Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR), 2024), featuring an accelerated and more robust algorithm as well as a broader empirical evaluation. The code and data relative to this article are available at https://github.com/RobinBruneau/RNb-NeuS2.
@@ 1 Introduction
The recovery of fine-grained structures is a long-standing bottleneck in 3D surface reconstruction. Traditional multi-view stereo (MVS) methods [26, 75] often produce overly smoothed surfaces, struggling with sharp edges and intricate geometries [76, 27]. Recent neural approaches have advanced surface reconstruction in this direction by shifting from local, patch-based optimisation to global, pixel-wise optimisation [97, 98, 66, 85]. Notably, methods such as HF-NeuS [87], PET-NeuS [89], NeuS2 [88], and Neuralangelo [54] excel in capturing fine surface details – at least, when the density of viewpoints is high [7, 58].
@@
This paper builds upon and significantly extends our previous conference publication [7], where the foundational concepts of this two-stage approach were introduced. Therein, we presented the key idea to re-parametrise normal and reflectance priors into simulated radiance values on a per-pixel basis, facilitating seamless incorporation into existing surface reconstruction pipelines, notably neural volume rendering (NVR) frameworks based on signed distance functions (SDF). This innovative way to perform multi-view normal (and reflectance) integration yields a state-of-the-art multi-view photometric stereo (MVPS) technique, summarised in Figure 1, which we evaluated on the DiLiGenT-MV dataset [53] using the SDM-UniPS method [38] for normal and reflectance estimation.
Figure 1: Overview of the surface reconstruction pipeline “RNb-NeuS” proposed in [7]. From multi-view multi-light data, photometric stereo (noted PS) first estimates per-view normal maps $\{\mathbf{n}_{k}\}$ and optionally reflectance maps $\{\mathbf{r}_{k}\}$. These estimates are re-parametrised into radiance vectors $\mathbf{v}_{k}=\mathsf{F}(\mathbf{n}_{k},\mathbf{r}_{k},\mathsf{L}_{k})$ via a physically-based rendering function $\mathsf{F}$ under simulated lightings $\mathsf{L}_{k}$. An implicit neural representation, consisting of a signed distance function $f$ and a reflectance $\boldsymbol{\rho}$, is then learned by minimising the discrepancy between $\mathbf{v}_{k}$ and its volume-rendered counterpart $\tilde{\mathbf{v}}_{k}(f,\boldsymbol{\rho})$. The final surface mesh is extracted as the zero-level set of $f$ after optimisation.
@@
[…]
@@
  • In addition to volume rendering, we validate our re-parametrisation in a traditional patch-based MVS framework, demonstrating both its broad applicability and its ability to achieve exact integration in a noiseless setup (see Section 4).
@@
The rest of this article is organised as follows. We first review the relevant literature in Section 2. Then, we show in Section 3 how to jointly re-parametrise normal and reflectance data into simulated radiance values. We embed this re-parametrisation in a traditional MVS framework in Section 4, to demonstrate the feasibility of exact surface reconstruction from normals and reflectance cues. To ensure robustness, in Section 5 we then turn our attention to replacing this framework by a volume rendering one based on neural implicit surfaces. Extensive experiments validating the proposed approach are then conducted in Section 6, before our conclusions are drawn in Section 7.
@@ 2.1 Normal Estimation
Single-image normal estimation is highly beneficial for 3D scene reconstruction, since it preserves local geometry without metric ambiguity [101, 84]. Recent approaches predominantly leverage learning-based methodologies [51, 86, 69, 70, 56, 20, 4, 95]. However, direct training on real data remains challenging, as normal labels cannot be directly captured by current sensor technology. Typically, normals are thus inferred from depth maps [79, 23], but inaccuracies persist despite various correction strategies [3, 60]. Therefore, the most recent works employed large-scale datasets of synthesised data such as OmniData [22, 44], along with smart architectural choices, e.g. per-pixel ray direction modelling in DSINE [2] or ViTs and diffusion models in StableNormal [99].
Photometric stereo [91], on the other hand, is also a monocular technique yet it analyses multiple images captured under varying lighting. The historical inverse problem-based PS approach, which fits normals and reflectance to a physics-based image formation model, has been employed in the calibrated (known illumination) scenario [28, 92, 77, 40, 39], in the uncalibrated one [33, 11, 1, 25, 72], as well as under unknown and spatially-varying lighting [5, 71, 63, 31, 30]. However, switching from the inverse problem framework to the deep learning paradigm has recently revolutionised PS. Limitations of traditional physics-based models, especially regarding non-Lambertian effects, have thereby been addressed for both calibrated  [74, 35, 12, 36, 42, 90] and uncalibrated PS [78, 80, 13, 14, 45, 52], and the most modern techniques make lighting context implicit, which both improves performances [43] and allows tackling “universal” (unknown and spatially-varying illumination) PS [37, 38, 32, 15].
In multi-view single-light contexts, traditional methods, such as Colmap’s multi-view stereo [75], infer normal maps by analysing patch orientations [6]. Differently, [9] estimate normals from image grey level variations and use them in a slanted plane-sweeping algorithm. An alternative paradigm computes normals post-MVS by deriving them from depth maps, as demonstrated in RegSDF [102], which utilises depth from Vis-MVSNet [103]. This depth-based normal estimation is MVS-method agnostic, applicable to both traditional [26, 93] and deep learning-driven methods [96, 29, 103].
@@ 2.2 Reflectance Estimation
In multi-view frameworks, methods such as Ref-NeRF [81] extend neural radiance fields (NeRF) to explicitly model reflections, improving the reconstruction of specular and glossy surfaces. Techniques like SHINOBI [24] concurrently optimise geometry, reflectance, and illumination, enabling detailed relightable 3D assets. More recently, DiffusionRenderer [55] provided a unified solution for high-quality normal and reflectance reconstruction from video sequences, by bridging monocular and multi-view estimation through diffusion models for neural inverse rendering.
@@ 2.3 Multi-view Normal Integration
In a multi-view single-light setting, MonoSDF [101] combined a normal and depth consistency loss derived from monocular predictions [22] with a radiance constraint, a common approach in NeRF-like [62] methods. Gaussian Surfels [18] uses monocular normal predictions to constrain 3D Gaussian Splatting (3DGS) [48], while
@@
Yet, multi-view normal integration is much more studied under the prism of multi-view photometric stereo. This problem was first addressed by [34], through the optimisation of a loss combining a rendering term and a discrepancy between photometric stereo normals and the optimised mesh. Note that this first approach required neither prior knowledge of camera poses nor illumination conditions. Similarly, [67, 68] simultaneously estimated normals, reflectance, and illumination through uncalibrated PS, leveraging the 3D mesh normals to resolve ambiguities and refining surface details. [57] later formulated the problem within an SDF representation, achieving superior surface detail reconstruction compared to [68]. Further refinements were made by [53], who enhanced a 3D mesh by propagating SfM points following the method of [65], and validating their method through the introduction of the publicly available dataset “DiLiGenT-MV”.
@@
Neural surface reconstruction methods have recently emerged as a promising alternative, by enforcing alignment between per-view normal maps and the gradient of a neural SDF. [46] for instance constrained the SDF optimisation by CNN-PS normals  [35] and MVS depths [83], incorporating uncertainty measures to mitigate conflicts in predictions. [47] further added a neural volume rendering loss, improving robustness to various material types. This resulted in a multi-objective optimisation comprising three loss terms. However, as in [46], the reliance on uncertainty-based hyperparameter tuning did not fully resolve conflicts between objectives, leading to potential loss of fine-grained details. PS-NeRF [94] introduced a two-stage solution, where the first stage achieves multi-view normal integration by aligning surface gradients with SDPS-Net normals [13], and the second stage leverages UNISURF [66] to optimise geometry, reflectance and illumination, modelled by multi-layer perceptrons (MLPs). However, its reliance on the directional illumination assumption limited its generalizability. NPL-MVPS [59] relaxed the directional illumination assumption by considering a near-light model, initially enforcing alignment between SDF gradients and UniMS-PS normals [32]. It then jointly optimises shape and reflectance via a rendering loss that explicitly models light attenuation and cast shadows. Most recently, SuperNormal [10] adopted a similar initialisation approach to PS-NeRF and NPL-MVPS, utilizing SDM-UniPS normals [38], but significantly improved efficiency with multi-resolution hash encoding and directional finite differences, achieving nearly double the training speed.
@@
In contrast to these prior methods, the novel approach we propose in the rest of this article formulates multi-view normal integration as a single-objective optimisation problem, through joint re-parametrisation of normals and reflectance. Let us now introduce this re-parametrisation.
3 Normal and Reflectance Re-parametrisation
@@
Surface normals and reflectance traditionally represent distinct types of information: normals describe the geometric orientation of surfaces, while reflectance captures intrinsic photometric properties of surface materials. This inherent heterogeneity complicates their simultaneous optimisation, typically necessitating separate processing steps, e.g., depth triangulation from multi-view data followed by the fusion of this depth information with normal maps using multi-objective optimisation frameworks [46, 47].
@@
To address this issue, we introduce a re-parametrisation approach designed to unify surface normals and reflectance into homogeneous quantities, by mapping both of them to simulated radiance values under varying illumination conditions. This radiance-based parametrisation enables a unified optimisation reducing the need for additional regularisation, commonly required in multi-objective contexts [101, 102], consequently enhancing consistency and computational efficiency. Indeed, a key motivation for radiance-based re-parametrisation is its compatibility with existing photometric cost-minimisation frameworks employed in classical multi-view stereo [26, 75] and neural volume rendering methods [85, 98, 54]. This incorporation facilitates optimisation of photometric consistency across viewpoints or between input and rendered images.
@@ 3.1 Input Data
Given a set of $N$ viewpoints, we assume the availability of corresponding normal maps $\mathsf{N}_{i}$ and reflectance maps $\mathsf{R}_{i}$, indexed by $i\in\{1,\dots,N\}$. Each viewpoint map comprises $m$ pixels, indexed by $k\in\{1,\dots,m\}$, such that:
$\mathsf{N}_{i}=\{\mathbf{n}_{i,k}\}_{k\in\{1,\dots,m\}},\quad\mathsf{R}_{i}=\{\mathbf{r}_{i,k}\}_{k\in\{1,\dots,m\}}.$ (1)
Therein, the outward normal for the $k$-th pixel in the $i$-th view is given by a unit vector $\mathbf{n}_{i,k}\in\mathbb{S}^{2}$, and the reflectance by a parameter vector $\mathbf{r}_{i,k}\in\mathbb{R}^{q}$ with dimension $q$ corresponding to the specific reflectance model used. The former is expressed, using the known camera poses, in world coordinates. If reflectance data is unavailable, simply setting $\mathbf{r}_{i,k}=[1]~\forall(i,k)$ allows our framework to be used exclusively for multi-view normal integration.
3.2 Re-parametrisation
Our re-parametrisation approach transforms each couple of normal and reflectance vectors $(\mathbf{n}_{i,k},\mathbf{r}_{i,k})$ into a vector $\mathbf{v}_{i,k}$ of homogeneous radiance values, simulated using a physically-based rendering (PBR) model:
$\mathbf{v}_{i,k}=\mathsf{F}(\mathbf{n}_{i,k},\mathbf{r}_{i,k},\mathsf{L}_{i,k}),$ (2)
with $\mathsf{L}_{i,k}\in\mathbb{R}^{n\times l}$ the illumination conditions (chosen specifically for the $k$-th photosite of the $i$-th view) and $\mathsf{F}:\,\mathbb{S}^{2}\times\mathbb{R}^{q}\times\mathbb{R}^{n\times l}\to\mathbb{R}^{n\times q}$ the PBR function (for simplicity, the same model is used for all pixels). Here, $l$ stands for the dimensionality of the light and $n$ for that of the radiance representation, and both of them depend on the choice of a particular model.